Martha has 2 more dimes than quarters. If she has $3.35 altogether, how many of each coin does she have?
Found 2 solutions by VFBundy, greenestamps:
Answer by VFBundy(438)

# of quarters = x
# of dimes = x + 2

0.25x + 0.10(x + 2) = 3.35

0.25x + 0.10x + 0.20 = 3.35

0.35x + 0.20 = 3.35

0.35x = 3.15

x = 9

# of quarters = x = 9
# of dimes = x + 2 = 11

Check it:

"Martha has 2 more dimes than quarters." 9 quarters and 11 dimes...so it checks.

"If she has $3.35 altogether..." 9 quarters @ $0.25 each equals $2.25 and 11 dimes @ $0.10 each equals $1.10. $2.25 plus $1.10 equals $3.35...so it checks.

Answer by greenestamps(13208)



(1) Count the 2 "extra" dimes and set them aside. What is left is equal numbers of dimes and quarters, with a total value of $3.35-$0.20 = $3.15.

(2) The value of 1 quarter and 1 dime is $0.35; the number of quarters and dimes it takes to make $3.15 is $3.15/$0.35 = 9.

(3) So the total number of quarters is 9, and the total number of dimes is 9+2 = 11.